Prove that for all set A, B, and C, (A \ C) ∩ (B \ C) ∩ (A \ B) = ∅.
Answer #1

A-C B-C A-B (A-C)B-Cn(A-B) From venn diagram it is observed that there are no common portion(A/C)cap (B/C)cap (A/B)=(Acap C^c)cap(Bcap C^c)cap(Acap B^c)\ =(Acap C^c)cap ((Acap C^c)cap (Bcap B^c))=(Acap C^c)cap (Bcap B^c)=(Acap C^c)cap phi\ =phi